5)If the recessive allele, a, has a frequency of 0.18 how many AA individuals are there in a randomly mating population of 377 individuals? (rounded up to the nearest whole individual)

6)  The Himalayan colour pattern in rabbits is controlled by a single locus with three alleles; F for full colour, H for Himlalayan and A for albino. The F allele is dominant over the H and A alleles while the H allele is dominant over the A allele. If the number of individuals with the Himalayan colour pattern represent 36% of the population and the albinos represent 11% of the population, what is the frequency of the F allele? (rounded to 3 decimal places)
Answer #1

5. The answer is

According to Hardy Weinberg's law

p+q = (p+q)^2= p^2+2pq+q^2 = 1.

Where
p represents dominant allele i.e A and
q represents recessieve allele i.e. a.

The proportion of recesseive allele = 0.18, so, the proportion of dominant allele = 1-0.18 = 0.82.

So, the proportion of AA = 0.82*0.82 = 0.6724

For 377 individuals the AA frequency = 0.6724*377 = 253.4948 ~ 253.

ACCORDING TO CHEGG GUIDELINES WE HAVE TO ANSWER ONE QUESTION AT A TIME. POST THE REST AS SPERATE QUESTIONS, THEN I CAN HELP YOU.